# Entity-aware Cross-lingual Claim Detection for Automated Fact-checking

### v1

@@ Appendix A Statistics
Appendix B Analysis
Attention Weights for Sample Sentences
@@ Appendix B Analysis
It can be observed, that EX-Claim attends to important words in the sentences, whereas X-Claim generates attention weights distributed across most of the words. Further, an interesting pattern of named entities attending to themselves or other named entities more compared to other words can be observed across all three sentences.
Correlation between Confusion Matrix and Selected Features
@@
Similarly, the number of disease names appearing in a claim is positively correlated with a false positive rate. This could be the COVID-19 discussion in the training data making the model to classify the claims with disease names as verifiable. Further, the number of media names is positively correlated with false negative rate as discussed in Section Error Analysis. This suggests that references to media, such as images and videos, can either enhance or hinder verifiability, depending on the context and reach.
(a) S10 : EX-Claim-CP
(b) S10 : X-Claim
(c) S11 : EX-Claim-F
@@
Appendix C Implementation
We fine-tuned the mBERT [^5] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-R [^6] https://huggingface.co/FacebookAI/xlm-roberta-base, and mT5 models published in HuggingFace to obtain the transformer-based baseline models. Especially, mT5 is available in various sizes. We used mT5-large [^7] https://huggingface.co/google/mt5-large as this model was the largest model among the variations which we could accommodate with the same resources utilized for training and testing other models. Similarly, the open-source large language models (LLMs) available in HuggingFace were used to obtain the LLM baselines, Llama2888https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral 7B [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, and Phi3101010https://huggingface.co/microsoft/Phi-3-mini-4k-instruct. We used the same experimental setting of (Li, Panchendrarajan, and Zubiaga 2024). The prompt 1 used to fine-tune LLMs is a modified version of (Li, Panchendrarajan, and Zubiaga 2024).
@@ Appendix C Implementation
through fact-checking. If the statement
presents assertions, facts, or claims,
respond with ’Yes’. If the statement is
purely opinion-based, trivial, or does not
contain any verifiable information or claims,
respond with ’No’.
@@
All the experiments were conducted using Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT (King, Butcher, and Zalewski 2017). Specifically, 1 GPU (Volta V100 or Ampere A100) with 8 CPU cores, each composed of 11 GB memory was used to train and test all the models. All the models were fine-tuned by extracting the vector representation of the special token from the final layer and applying a fully connected layer on it for classification. We used only the encoder of the mT5 model for obtaining vector representation of input words. For Named Entity Recognition (NER), we used the WikiNEuRal  [^11] https://huggingface.co/Babelscape/wikineural-multilingual-ner model published in HuggingFace and the resources published for MultiNERD [^12] https://github.com/Babelscape/multinerd/tree/master for fine-tuning the fine-grained NER model. Similarly, the mGENRE [^13] https://huggingface.co/facebook/mgenre-wiki model published in HuggingFace was used for Entity Linking (EL). License of the models used in the experiments can be found at the links provided.

### v3

@@ A.3.2 Kazemi Dataset
Appendix B Results
Baseline Entity Type Categories
Language
mBERT
XLM-R
mT5
Llama2
Mistral
Phi3
X-Claim
EXN-Claim
EXP-Claim
Arabic 0.692 0.654 0.677 0.752 0.736 0.718 0.731 0.733 0.735
Bengali 0.725 0.655 0.556 0.697 0.642 0.552 0.713 0.748 0.764
Bulgarian 0.783 0.633 0.729 0.835 0.815 0.812 0.782 0.79 0.796
Czech 0.707 0.647 0.72 0.808 0.786 0.755 0.741 0.774 0.769
Dutch 0.68 0.698 0.634 0.72 0.657 0.663 0.716 0.717 0.718
English (Kaz) 0.738 0.769 0.748 0.764 0.75 0.719 0.755 0.757 0.755
English (CT) 0.724 0.658 0.688 0.769 0.777 0.755 0.75 0.762 0.761
French 0.638 0.698 0.658 0.736 0.739 0.738 0.703 0.722 0.725
German 0.751 0.714 0.722 0.808 0.822 0.816 0.789 0.778 0.775
Greek 0.677 0.64 0.71 0.697 0.687 0.679 0.735 0.747 0.759
Gujarati 0.661 0.669 0.649 0.457 0.582 0.455 0.77 0.77 0.751
Hindi 0.631 0.695 0.613 0.792 0.809 0.6 0.766 0.767 0.769
Hungarian 0.667 0.651 0.665 0.757 0.727 0.637 0.68 0.683 0.682
Italian 0.721 0.647 0.704 0.764 0.745 0.736 0.709 0.731 0.723
Japanese 0.716 0.658 0.693 0.759 0.771 0.74 0.715 0.732 0.746
Kannada 0.535 0.567 0.579 0.371 0.459 0.342 0.628 0.674 0.676
Korean 0.691 0.635 0.623 0.69 0.72 0.653 0.712 0.718 0.709
Marathi 0.588 0.64 0.603 0.643 0.532 0.46 0.668 0.666 0.664
Malayalam 0.79 0.564 0.685 0.374 0.151 0.141 0.845 0.84 0.835
Persian 0.673 0.669 0.622 0.713 0.667 0.602 0.704 0.716 0.718
Russian 0.709 0.667 0.641 0.766 0.764 0.734 0.715 0.724 0.724
Spanish 0.738 0.693 0.683 0.779 0.735 0.77 0.787 0.783 0.784
Swedish 0.707 0.688 0.654 0.772 0.779 0.76 0.777 0.778 0.787
Tamil 0.623 0.744 0.789 0.439 0.398 0.264 0.864 0.866 0.861
Thai 0.628 0.648 0.644 0.716 0.698 0.673 0.726 0.716 0.715
Turkish 0.699 0.648 0.664 0.775 0.82 0.745 0.748 0.759 0.758
Ukrainian 0.688 0.597 0.604 0.764 0.724 0.719 0.644 0.688 0.707
Vietnamese 0.727 0.672 0.676 0.735 0.694 0.689 0.715 0.717 0.729
Table 8: F1-Score of the Claim Detection Models at Language-level
B.1 Language-level Performance
Table 8 lists the language-level F1-score of the claim detection models across 27 languages present in the test datasets.
B.2 Machine Translation Performance
Since the synthetic dataset was generated using machine translation, we assess the quality of the translation model—Microsoft Azure AI Translator Junczys-Dowmunt (2019). For this evaluation, we utilize FLORES-200 Team et al. (2022), a multilingual benchmark dataset that supports English-to-200+ language translation. To approximate the conditions of our synthetic dataset, we sampled sentence pairs from the five source languages represented in the CheckThat! training set to 18 target languages present in the synthetic dataset. For each target language, we selected 250 translation pairs, resulting in a total of 4,500 sentence pairs. This sampling strategy is designed to mimic the structure of our synthetic dataset, where each target language includes 250 translated claims.
Metric Score
BLEU 26.23
CHRF 54.35
CHRF++ 51.63
Table 9: Machine Translation Performance of Microsoft Azure AI translator
Table 9 presents the performance of the machine translation model on the sampled translation pairs, evaluated using multiple translation quality metrics. The model achieves competitive scores on the character-based metrics CHRF and CHRF++, with the CHRF++ score notably surpassing the baseline results reported in the original FLORES-200 dataset paper Team et al. (2022).
Appendix C Analysis
C.1 Entity-type Embedding
Figure 7: Two-dimensional embedding visualization of EXN-Claim Model
Unlike the cross-lingual word embedding which is directly obtained from the pre-trained model, EX-Claim learns a dense vector representation for each entity type. Figure 7 shows the 2D visualization of the entity type embeddings learned by the EXN-Claim model. The 2D vector representation of the embeddings is obtained using the t-SNE [^4] https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html dimension reduction technique. Further, the embedding representations were clustered using the KMeans clustering technique (Sinaga and Yang, 2020), and the number of optimal clusters was chosen using the Silhouette score (Rousseeuw, 1987).
The EXN-Claim model learns various groups of entity types. Notably, this model clusters Animal, Disease, Person, Biological Entity, Disease, Plan and Mythological Entity into a single group, likely influenced by the COVID-19 discussions present in the training data. This observation highlights the model’s tendency to learn context-specific entity-type embeddings during training.
C.2 Attention Weights for Sample Sentences
@@
It can be observed that EX-Claim attends to important words in the sentences, whereas X-Claim generates attention weights distributed across most of the words. Further, an interesting pattern of named entities attending to themselves or other named entities more compared to other words can be observed across all three sentences.
C.3 Ablation Study
Model F1-Score
X-Claim 0.735
X-Claim - NPL 0.688
X-Claim - NAL 0.714
Table 10: Ablation Results of X-Claim Model
Model F1-Score
EXN-Claim 0.745
EXN-Claim - NEL 0.733
EXP-Claim 0.746
EXP-Claim- - NEL 0.733
Table 11: Ablation Results of EX-Claim Models
C.4 Correlation between Confusion Matrix and Selected Features
@@
Similarly, the number of disease names appearing in a claim is positively correlated with a false positive rate. This could be the COVID-19 discussion in the training data making the model to classify the claims with disease names as verifiable. Further, the number of media names is positively correlated with both true negative and false negative rates as discussed in Section 6.3. This suggests that references to media, such as images and videos, can either enhance or hinder verifiability, depending on the context and reach.
(a) S4 : EXN-Claim
(b) S4 : X-Claim
(c) S11 : EXN-Claim
@@
Appendix D Implementation
Model # Parameters Multilingual? Training time
Transformer-based Baselines mBERT 110M ✓ Yes (104 langs) 20 sec
XLM-R 270M ✓ Yes (100 langs) 20 sec
mT5 Large Encoder 600M ✓ Yes (101 langs) 20 sec
LLMs LLaMA 2 7B Limited 3 hours & 16 mins
Mistral 7B Limited 3 hours & 25 mins
Phi-3 3.8B Limited 2 hours & 12 mins
Entity-aware Models X-Claim 197K ✓ Yes (100 langs) 20 sec
EXN-Claim 201K
✓ Yes (100 langs
+ NER model langs)
20 sec
EXP-Claim 206K
✓ Yes (100 langs
+ NER model langs
& EL model langs)
21 sec
Table 12: Comparison of Model Parameters, Multilingual Support, and Training Time
D.1 Resources
All the experiments were conducted using Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT King et al. (2017). Specifically, 1 GPU (Volta V100 or Ampere A100) with 8 CPU cores, each composed of 11 GB memory was used to train and test all the models.
D.2 Models
We fine-tuned the mBERT [^5] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-R [^6] https://huggingface.co/FacebookAI/xlm-roberta-base, and mT5 models published in HuggingFace to obtain the transformer-based baseline models. Especially, mT5 is available in various sizes. We used mT5-large [^7] https://huggingface.co/google/mt5-large as this model was the largest model among the variations which we could accommodate with the same resources utilized for training and testing other models. Similarly, the open-source large language models (LLMs) available in HuggingFace were used to obtain the LLM baselines, Llama2888https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral 7B [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, and Phi3101010https://huggingface.co/microsoft/Phi-3-mini-4k-instruct. We used the same experimental setting of Li et al. (2024). The prompt 1 used to fine-tune LLMs is a modified version of Li et al. (2024).
@@ D.2 Models
through fact-checking. If the statement presents
assertions, facts, or claims, respond with ’Yes’.
If the statement is purely opinion-based, trivial,
or does not contain any verifiable information or
claims, respond with ’No’.
@@
All the models were fine-tuned by extracting the vector representation of the special token from the final layer and applying a fully connected layer on it for classification. We used only the encoder of the mT5 model for obtaining vector representation of input words. For Named Entity Recognition (NER), we used the resources published for MultiNERD [^11] https://github.com/Babelscape/multinerd/tree/master for fine-tuning the fine-grained NER model. Similarly, the mGENRE [^12] https://huggingface.co/facebook/mgenre-wiki model published in HuggingFace was used for Entity Linking (EL).
D.3 Model Comparison
Table 12 compares claim detection models based on the number of parameters, multilingual support, and training time. Both EXN-Claim and EXP-Claim incur an additional average latency of 5.2 ms per sentence for named entity recognition. Furthermore, EXP-Claim requires an average of 0.15 seconds per entity to perform entity linking.
D.4 Hyperparameters
Hyperparameter Value
Projection Size 256
Learning Rate 3e-5
Entity Embedding Size 128 (X, EX), 256 (EXP)
Entity Linking Threshold -0.15
Table 13: Hyperparameters Used
Table 13 lists the hyperparameters used. Following Devlin et al. (2019), we set the maximum length of the input sequence to 128, and use Adam as the optimizer. We train the models for 30 epochs and choose the model with the lowest validation loss as the best model.